OpenStudy (anonymous):

Define the following functions . f(x)=x^2; g(x)=1/√x a.f+g b.f-g c.f⋅g d.f/g e.g/f

OpenStudy (gorv):

.f+g=x^2+(1/√x) f-g=x^2-(1/√x) f⋅g=x^2*x^(-1/2)=x^3/2 f/g=x^2*x^(1/2)=x^5/2 g/f=(1/√x)/x^2=x^-1/2 * x^-2=x^-5/2

OpenStudy (anonymous):

a&b cant be operated?

OpenStudy (anonymous):

@MysteryOwl A=x not =0 b=x not=0 c=for all value of "x" d=for all value of 'x' e= not =0

OpenStudy (anonymous):

uhmm.. can you explain further please?

OpenStudy (anonymous):

@MysteryOwl i mean 'x'can take any real value ,except those which are mentioned earlier by me. ok!

OpenStudy (anonymous):

@metrxqin hello

OpenStudy (anonymous):

hello,i'm just looking around

OpenStudy (anonymous):

its ok :))

OpenStudy (cggurumanjunath):